OpenStudy (anonymous):

How do you graph the function f(x)=13x+900

Nnesha (nnesha):

start with y-intercept y=mx+b slope intercept form where m is slope b is y-intercept and slope = rise over run

Nnesha (nnesha):

b is what in that equation ?

OpenStudy (anonymous):

B is 900

Nnesha (nnesha):

yes right so draw a point at 900  now slope is 13 which is same as \[\rm \frac{ 13 }{ 1 }=\frac{ rise }{ runn }\] go up 13 units and then right one